OpenStudy (anonymous):

please help. (: Which factor can you cancel from the numerator and denominator to simplify 7 times x minus 21 all over 18 minus 6 times x.? [Hint: (a−x) = − (x − a)] 6 over 7. x − 3 7 over 6. x + 3

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

can you write that in numerals?

OpenStudy (johnweldon1993):

So lets see \[\large \frac{7x - 21}{-6x + 18}\] right?

OpenStudy (johnweldon1993):

What are both terms on the top divisible by? What are both terms on the bottom divisible by?

OpenStudy (anonymous):

\[\frac{ 7x-21 }{18--6x }\]

OpenStudy (johnweldon1993):

Oh its minus minus on the bottom?

OpenStudy (anonymous):

its minus at the bottom, yeah.

OpenStudy (johnweldon1993):

Okay wait, stick with me lol...because I see a double negative in the bottom that you wrote so is it \(\large \frac{7x - 21}{18 - (-6x)}\) or \(\large \frac{7x - 21}{18 - 6x} \)

OpenStudy (anonymous):

------>>

OpenStudy (anonymous):

so you need to find a common factor.

OpenStudy (johnweldon1993):

Lol okay thank you \[\large \frac{7x - 21}{18 - 6x}\] What number is each term on the top divisible by? What number is each term on the bottom divisible by?

OpenStudy (anonymous):

bottom is divisible by 3?

OpenStudy (anonymous):

?

OpenStudy (johnweldon1993):

Oh sorry, it never told me you responded! Umm, yes it is divisible by 3...but it is also divisible by 6 right? we just kinda want the largest number here What about the top?

OpenStudy (anonymous):

7

OpenStudy (johnweldon1993):

Right, so you know the distributive property right? The one that says something like *just an example* \[\large 3(x + 2) = 3x + 6\]

OpenStudy (anonymous):

yeah

OpenStudy (johnweldon1993):

Okay good, so you're with me if I write \[\large \frac{7(x - 3)}{6(3 - x)}\]

OpenStudy (anonymous):

yep .;p

OpenStudy (johnweldon1993):

haha well good ;p So looking at that hint you got in the question \[\large (a - x) = -(x - a)\] We can now write this as \[\large \frac{7(x - 3)}{-6(x - 3)}\]

OpenStudy (anonymous):

okay

OpenStudy (johnweldon1993):

Make sense? Trying to do it so you follow along :D So now what cancels?

OpenStudy (anonymous):

\[\frac{ 7 }{ -6 }\]

OpenStudy (johnweldon1993):

Well that's what's left after the \(\large (x - 3)\) cancels ;P

OpenStudy (anonymous):

Lol. Thank you!

OpenStudy (johnweldon1993):

Not a problem :D